OpenStudy (anonymous):

Convert the polar equation to rectangular form and identify the graph. r= 7/ (2cosΘ+5sinΘ)

OpenStudy (anonymous):



OpenStudy (anonymous):

y = r sin(theta) x = r cos(theta) x^2 + y^2 = r^2 plug and chug

OpenStudy (anonymous):

yea so how would i plug it in

OpenStudy (anonymous):

cuz i did this last week and i kinda forgot sorry

myininaya (myininaya):

Try getting your cos and sin with r over there by multiplying both sides by 2cos(theta)+5sin(theta)

OpenStudy (anonymous):

okay. i did that

myininaya (myininaya):

So you have: \[2rcos(\theta)+5rsin(\theta)=7 ?\]

OpenStudy (anonymous):

why would you put r w/ each one of the terms isnt there just 1?

myininaya (myininaya):

\[r=\frac{7}{2\cos(\theta)+5 \sin(\theta)} => r(2\cos(\theta)+5 \sin(\theta))=7\] Distributive property!

OpenStudy (anonymous):

oh ok i gotchu

myininaya (myininaya):

you should have it from here. Use the equations sour gave you.

OpenStudy (anonymous):

thannk you(:

OpenStudy (anonymous):

what if the question's like r= -3sinΘ ?

myininaya (myininaya):

Try multiplying r on both sides recall rsin(theta)=y and r^2=x^2+y^2

OpenStudy (anonymous):

hmm ok...

OpenStudy (anonymous):

so r^2= -3rsinΘ?

myininaya (myininaya):

right so now you are able to replace rsin(theta) with y and r^2 with x^2+y^2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

then youd have?

OpenStudy (anonymous):

the square root of that right?

myininaya (myininaya):

just do exactly what I said replace r sin(theta) with y replace r^2 with x^2+y^2

OpenStudy (anonymous):

i did tht

myininaya (myininaya):

you are done what else do you want?

OpenStudy (anonymous):

is tht my final answer

OpenStudy (anonymous):

oh ok

myininaya (myininaya):

putting an equation in terms of just x and y ( or just x or just y) is called a cartesian equation if that was the directions to just right as a cartesian equation then you are done

myininaya (myininaya):

write*

OpenStudy (anonymous):

right, sorry. i missed a problem on my homework it was r= 2sinΘ-4cosΘ

myininaya (myininaya):

multiply both sides by r

OpenStudy (anonymous):

do you square both sides 1st

OpenStudy (anonymous):

thts what i emant

myininaya (myininaya):

replace r^2 with x^2+y^2 replace r sin(theta) with y replace r cos(theta) with x

OpenStudy (anonymous):

ok i did that: x^2+y^2= 2y-4x

myininaya (myininaya):

yep. :) that is right you wrote the poloar equation as a cartesian equation if that is all the directions said to do then you are done.

OpenStudy (anonymous):

do you stop there

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

aand i missed r= 3+3cosΘ

OpenStudy (anonymous):

so i multiplied both sides by r 1st

OpenStudy (anonymous):

and got x^2+y^2=3r +3x but dont i need to get rid of the r?

myininaya (myininaya):

you could subtract both sides by 3x giving you \[x^2+y^2-3x=3r \] then square both sides and replace the r^2 with x^2+y^2

OpenStudy (anonymous):

why square both sides?

myininaya (myininaya):

I just prefer my answer without radicals and me squaring both sides will allow me to write my answer without radicals

OpenStudy (anonymous):

so what would be final answer then

myininaya (myininaya):

what if you with what I was saying take my equation and then square both sides and then replace the r^2 with x^2+y^2 you will have your final answer

OpenStudy (anonymous):

yeah but when you square the whole left side you get x^4+y^4-9x^2 = 3x^2+y^2?

myininaya (myininaya):

\[(x^2+y^2-3x)^2=(3r)^2 \] the us what you get when you square both sides your equation is not equivalent to that you squared each term sorta but not really it is whatever you do to one side of equation you can do to the other and the equation still holds

myininaya (myininaya):

\[\text {recall law of exponents} (3r)^2=3^2r^2=9r^2=9(x^2+y^2) \]

OpenStudy (anonymous):

so..

OpenStudy (anonymous):

im not following are you saying keep left side just squared in parentheses?

myininaya (myininaya):

It is easier than multiplying it out, don't you think?

OpenStudy (anonymous):

yes. so only change right side. like substitue

myininaya (myininaya):

You can multiply it out if you want. Just do it correctly, but the if the directions just say write as a rectangular (aka cartesian equation) than you are done. If the left hand side was already in terms of x and y. r^2 was the only thing that needed to be converted.